Exercise 1. prove the statement, Every rational number can be expressed as a quotient p/q so that p and q has no common factors. 


1
Expert's answer
2021-01-12T14:42:19-0500

A number is said to be rational number if it is expressed as "\frac{p}{q}" form, where "q\neq0" and "g.c.d" "(p,q)=1" .

Now to prove the statement two cases arises :

Case 1: If "p,q" have no common factor then "g.c.d" "(p,q)" =1.which is obvious.

Case 2: If "p,q" have common factor.

Let us take "r" is a common factor of "p" and "q" .

Then "p=r.m" for some "m\in Z"

"q= r.n" for some "n\in Z-"{0} [ since "q\neq0," then "n\neq0" ]

and "g.c.d" "(m,n)=1"

In this case "\frac{p}{q}" can be written in the form "\frac{m}{n}", where "g.c.d" "(m,n)=1" . Which is also in the form of "\frac{p}{q}" .

Therefore from these two cases we can conclude that every rational number can be expressed as a quotient of "\frac{p}{q}" , so that "p" and "q" have no common factors.